OpenStudy (anonymous):

Need help with f(x) (I'm working it out) The question is: given f(x) = x - 7 and g(x) = x^2, find: g(f(4)) I know that I have to work from the inside and find f(x) first, but wouldn't that be f(4)? f(x) = x - 7 would be 4 - 7 = -3 And then to find g(x) I have g(x^2)= x^2 Finally, I have x^2(3) I know I did this wrong but I'm not sure where I messed up.

OpenStudy (agent0smith):

g(f(4)), when you know f(4) = -3 becomes: g(-3) so find that.

OpenStudy (agent0smith):

If you know how to find f(4), you know how to to find g(-3)... plug in -3 in place of x.

OpenStudy (anonymous):

Okay, so I would put g(-3) = 3^2 = 9?

OpenStudy (agent0smith):

Yes.

OpenStudy (anonymous):

9 wouldn't be the final answer would it?

OpenStudy (agent0smith):

g(f(4)) = g(-3) = 9

OpenStudy (anonymous):

Or would I plug it back in like: 9(-3) I just end the problem when I find g?

OpenStudy (agent0smith):

Look at the steps: g(f(4)) = g(-3) = 9

OpenStudy (agent0smith):

Do you see any more possible steps to take?

OpenStudy (agent0smith):

Don't over-complicate things.

OpenStudy (anonymous):

Okay, I think I got it. I am doing online homework and it's asking me for the opposite now: f(g(4)) so I think I was just making it harder than it was! Thank you!

OpenStudy (agent0smith):

f(g(4)) again start from the inside and work outwards.

OpenStudy (anonymous):

g(4) = 4^2 = 16 Substitute 16 into f and have 16 - 7 = 9 f(g(4)) = 9

OpenStudy (agent0smith):

Yes.

OpenStudy (anonymous):

Thanks for your help, I understand it a lot more now!